The specific gravity of an object is 0.957, whereas that of seawater is 1.025.

What percent of an object is above the surface of the water?
Answer #1

Solution)

Given,

Here, Density of object/ Density of water= Volume of displaced water/volume of object

Substitute values,

=0.957/1.025=0.933

Peecentage of object above water=(1-0.933)*100= 6.7% (Ans)

============